The ratio of girls to boy in the class is 2:3. If there are 40 students in the class, how many are girls?
Answer by TimothyLamb(4379)

---
x = boys
y = girls
---
y/x = 2/3
x + y = 40
---
put the system of linear equations into standard form:
---
y = (2/3)x
x + y = 40
---
( (2/3)x - y = 0 )3
x + y = 40
---
2x - 3y = 0
x + y = 40
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = boys = 24
y = girls = 16
---
system matrix
2 -3
1 1
---
inverse of system matrix
0.2 0.6
-0.2 0.4
---
determinant of system matrix = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php